OpenStudy (anonymous):

A pulley system has an efficiency of 75.2 percent. The acceleration of gravity is 9.81 m/s 2 . How much of the rope must be pulled in if a force of 549 N is needed to lift a 130 kg desk 2.01 m? Answer in units of m

OpenStudy (anonymous):

\[G=mg = 130 kg \times9.81 \frac{ m }{ s^2 } = 130 \times 9.81 N\] Formula used, since amount of work (energy needed) is same : \[Work = F \times s =\eta \times W \] \[130 \times9.81 \times 2.01 = 0.752 \times (549 N \times s) \, (Nm)\] Solve for distance of rope s: \[s = \frac{130 \times9.81 \times 2.01}{0.752 \times 549 N} = .. \, (Nm)\]

OpenStudy (anonymous):

Additions \[s=\frac{130×9.81×2.01}{0.752×549}m \approx 6.21 m\]

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

First the needed work to lift 130 kg up 2.01 meters is: \[W=Fs=mgs = 130 kg \times9.81 \frac{ m }{ s^2 } \times 2.01 m= 130 \times 9.81 \times 2.01 Nm\]

OpenStudy (anonymous):

There is some clutter so I thought about reposting the whole thing ... since I couldn't edit the post once posted ..